OpenStudy (anonymous):

find four consecutive odd integers such that three x the sum of the first and second is 35 less than the fourth.

OpenStudy (callisto):

Let x be the smallest number, the other 3 numbers are x+2, x+4, x+6 3 [x+(x+2)] = (x+6) -35 3( 2x +2) = x-29 5x = -35 x=-7

OpenStudy (anonymous):

Thankk tou

OpenStudy (callisto):

Welcome, do you understand how to do it?